### 05 Mar 2026 v1

# The Geometric Inductive Bias of Grokking: Bypassing Phase Transitions via Architectural Topology

Abstract
Mechanistic interpretability has largely focused on post-hoc analysis of trained networks. In this work, we instead adopt an interventional approach: we test mechanistic hypotheses a priori by modifying architectural topology and observing the resulting training dynamics. We study grokking—the delayed generalization phenomenon observed when Transformers are trained on cyclic modular addition ($\mathbb{Z}_{p}$)—and investigate whether specific architectural degrees of freedom contribute to the prolonged memorization phase.
We identify two independent structural factors present in standard Transformers: unbounded representational magnitude and data-dependent attention routing. First, we introduce a fully bounded spherical topology that enforces $L_{2}$ normalization throughout the residual stream and normalizes the unembedding matrix with a fixed temperature scale. This constraint removes magnitude-based degrees of freedom and reduces grokking onset time by more than $20\times$ without weight decay.
Second, we introduce a Uniform Attention Ablation that overrides data-dependent query–key routing with a fixed uniform distribution, reducing the attention layer to a Continuous Bag-of-Words (CBOW) aggregator. Despite eliminating adaptive routing, LayerNorm models with uniform attention achieve 100% generalization across all seeds and bypass the grokking delay.
To evaluate whether this acceleration reflects task-specific geometric alignment rather than a generic optimization stabilizer, we employ non-commutative $S_{5}$ permutation composition as a negative control. Enforcing the same spherical constraints on $S_{5}$ does not accelerate generalization, suggesting that eliminating the memorization phase depends strongly on aligning architectural priors with the intrinsic symmetries of the task. Together, these findings provide interventional evidence that architectural degrees of freedom substantially influence the presence and duration of grokking, suggesting a predictive architectural perspective on training dynamics.
Keywords: Grokking, Mechanistic Interpretability, Inductive Bias, Architectural Intervention, Transformer, Representational Geometry
1 Introduction
Neural networks often exhibit complex and unintuitive learning dynamics. One of the most striking examples is grokking: a delayed phase transition in which a model achieves near-perfect training accuracy while test accuracy remains low, followed by a sudden transition to full generalization after prolonged optimization. Originally observed in small models trained on algorithmic tasks, grokking has become a central phenomenon for studying the relationship between optimization, generalization, and implicit bias.
In parallel, mechanistic interpretability aims to uncover the internal algorithms learned by neural networks. For example, prior work has shown that Transformers trained on modular addition eventually implement structured representations related to discrete Fourier features (Nanda et al., 2023). However, much of this work is post-hoc: researchers analyze models only after grokking has occurred, inferring mechanisms from frozen weights.
In this work, we take a complementary interventional approach. Rather than analyzing a trained model, we modify architectural structure before training to test whether specific representational degrees of freedom influence grokking dynamics. Our central hypothesis is that standard Transformer architectures possess degrees of freedom that exceed the minimal symmetry requirements of commutative, periodic tasks such as modular addition. As demonstrated by Zhong et al. (2023), this excess capacity allows standard networks to converge on diverse, sometimes messy, piecewise solutions (the “Pizza” algorithm) rather than the elegant, continuous Fourier solution (the “Clock” algorithm). We hypothesize that these additional degrees of freedom enable memorization-heavy solution pathways that delay the emergence of invariant representations (Zhong et al., 2023; Lei and Xu, 2025; Prieto et al., 2025).
We isolate and study two independent architectural factors:
Path A: The Magnitude Degree of Freedom.
In standard Transformers, information can be encoded in both the direction and magnitude of residual stream vectors. Prior work has suggested that unconstrained magnitude growth can influence training dynamics in grokking settings (Prieto et al., 2025). To examine the role of this degree of freedom, we introduce a Fully Bounded Spherical Topology that enforces strict $L_{2}$ normalization throughout the residual stream and applies a normalized unembedding matrix with a fixed temperature scale. This intervention removes the model’s ability to encode information in vector norm and constrains logits to a bounded cosine geometry. Empirically, this modification substantially accelerates generalization, reducing grokking onset from $\sim$54,160 epochs in standard LayerNorm baselines to $\sim$2,100 epochs without weight decay.
Path B: The Routing Degree of Freedom.
Transformers also possess flexible, data-dependent attention routing through learned query-key interactions. However, recent theoretical results demonstrate that modular addition can be implemented using uniform token aggregation (Huang and Li, 2025). Motivated by this, we introduce a Uniform Attention Ablation that overrides learned query-key scores with a fixed uniform distribution, eliminating adaptive routing and reducing the attention mechanism to a Continuous Bag-of-Words (CBOW) aggregator. Despite removing routing flexibility, LayerNorm models with uniform attention achieve 100% peak test accuracy across 10 independent seeds and do not exhibit a prolonged grokking delay.
Summary of Contributions.
Through controlled architectural interventions, we show that (1) constraining representational magnitude and (2) removing data-dependent routing each independently and substantially reduce grokking delay in modular addition. These results demonstrate that grokking dynamics are highly sensitive to architectural degrees of freedom. Furthermore, using non-commutative $S_{5}$ permutation composition as a negative control, we observe that symmetry-aligned architectural constraints do not eliminate delayed generalization in tasks lacking commutativity, suggesting task-specific effects.
Overall, our findings support a predictive, geometry-aware approach to studying training dynamics, in which architectural priors are aligned with task symmetry to probe the mechanisms underlying delayed generalization.
2 Related Work
2.1 Grokking as Geometric Reorganization
Recent empirical studies suggest that the delayed generalization characteristic of grokking has been interpreted as a geometric phenomenon rather than a purely optimization-driven one. Observational work by Zheng et al. (2024) demonstrated that delayed generalization coincides with a rapid decrease in the effective radius and dimensionality of the representation manifold. They found that changes in this representational geometry are more tightly linked to the onset of generalization than standard optimization metrics.
Similarly, Lei and Xu (2025) characterized grokking as an asynchronous “construct-then-compress” dynamic. By tracking Jacobian transformations across local neighborhoods, they showed that standard networks spend the majority of their memorization phase building disjoint representations, only generalizing once these features are compressed into a globally coherent algorithmic geometry.
This geometric bottleneck extends to higher-order reasoning tasks as well. Minegishi et al. (2026) demonstrated that the emergence of analogical functors in Transformers relies heavily on structural alignment in the embedding space, which they quantified via the minimization of Dirichlet energy. Crucially, they noted that this alignment is highly sensitive to the presence of weight decay, reinforcing the necessity of regularization in forcing networks toward structured, low-energy geometric states.
Our work builds directly upon this observational foundation. While prior studies passively measure the slow emergence of geometric compression (Zheng et al., 2024; Lei and Xu, 2025), we treat geometry as an independent variable. By architecturally enforcing an $L_{2}$ spherical constraint, we structurally restrict the representational manifold from initialization, potentially reducing the need for an extended unconstrained construction phase.
This geometric bottleneck is intimately tied to the optimization objective itself. Recent theoretical work by Morwani et al. (2024) mathematically shows that the emergence of structured algebraic circuits—such as Fourier features for modular addition and irreducible representations for finite groups—is driven by the implicit margin-maximization bias of gradient descent under cross-entropy loss. In standard unconstrained networks, the prolonged grokking delay corresponds to the time required for this implicit bias to slowly sculpt the optimal mathematical geometry out of an unstructured parameter space. Our interventional methodology complements this analytical framework: by architecturally enforcing an $L_{2}$ spherical constraint, we explicitly restrict the representational manifold to the continuous geometry required by the target Fourier features, effectively short-circuiting the tedious margin-maximization process.
2.2 Periodic and Circular Inductive Biases
Parallel to observational studies on network geometry, recent work has explored explicitly embedding periodic or circular inductive biases into models to improve efficiency on algorithmic tasks. Huang and Li (2025) established a formal expressivity gap for modular addition, mathematically proving that a two-layer MLP with sinusoidal activations requires only a constant width of $2$ to exactly realize the task, whereas standard ReLU networks require a width that scales linearly with the sequence length or modulo. This demonstrates that matching the network’s non-linearity to the periodic nature of the target function drastically improves parameter efficiency and statistical learning guarantees.
Extending this principle to network initialization, Fernández-Hernández et al. (2025) introduced a deterministic sinusoidal weight initialization to replace standard stochastic methods. They demonstrated mathematically that random initialization inherently produces “skewed neurons” with imbalanced preactivations due to statistical fluctuations. By using structured harmonic functions, their initialization enforces a symmetric, zero-sum geometry from the first forward pass, preventing initial neuron skewness and accelerating early-stage convergence across various standard architectures.
These findings highlight that guiding a network toward a circular or symmetric geometry—whether through periodic activation functions (Huang and Li, 2025) or deterministic harmonic initialization (Fernández-Hernández et al., 2025)—can fundamentally alter optimization trajectories. Our methodology complements this intuition. Rather than altering the activation functions or the initialization weights, we impose a hard topological constraint (the $L_{2}$ hypersphere) directly on the residual stream, explicitly testing if architectural topology alone can effectively bypass the unconstrained memorization phase.
2.3 Fourier Features and Structural Continuity
Beyond localized activation functions, the structural topology of network layers plays a critical role in learning continuous and modular algorithms. Zhou et al. (2024) demonstrated that pre-trained large language models natively compute arithmetic by representing numbers via Fourier features. They observed a strict division of labor: MLP layers rely on low-frequency Fourier components to approximate magnitude, while attention layers utilize high-frequency components for precise modular classification. Crucially, Transformers trained from scratch fail to develop these Fourier features and consequently struggle with exact addition. This indicates that standard unconstrained architectures naturally lack the circular inductive bias required to efficiently solve these tasks without massive pre-training.
Addressing the unstructured nature of standard network components, Gillman et al. (2025) showed that standard linear classification heads frequently overfit to high-frequency noise when modeling continuous variables. To counteract this, they introduced the “Fourier head,” an architectural replacement that restricts the output distribution to a truncated Fourier series. By architecturally enforcing a smooth, continuous probability density, this topological constraint mathematically prevents the memorization of high-frequency noise, yielding substantial performance improvements in large-scale sequential decision-making and time-series forecasting.
Together, these works illustrate that unconstrained network components default to inefficient, high-frequency memorization strategies when tasked with continuous or periodic structures. While prior solutions involve relying on massive pre-training to discover Fourier features (Zhou et al., 2024) or explicitly constraining the output layer’s geometry (Gillman et al., 2025), our interventional approach tests whether enforcing a similar circular topology directly within the core residual stream can fundamentally alter the network’s optimization trajectory.
3 Formal Problem Setting and Methodology
To investigate the structural drivers of delayed generalization (grokking), we build upon the mechanistic interpretability framework established by Nanda et al. (2023). Their work demonstrated that transformers trained on modular addition construct continuous Fourier-based representations. In particular, learned attention heads organize token interactions such that downstream MLP layers compute trigonometric identities consistent with circular phase structure Zhou et al. (2024).
Standard transformer architectures, however, contain representational degrees of freedom that exceed the minimal mathematical requirements of commutative, periodic tasks. We identify two distinct excess degrees of freedom: unbounded vector magnitude in the residual stream, and asymmetric, data-dependent attention routing. We hypothesize that these additional degrees of freedom permit solution pathways that do not respect task symmetry, enabling symmetry-misaligned solution pathways that delay convergence to the structured Fourier circuit. Zheng et al. (2024); Lei and Xu (2025).
To evaluate this hypothesis, we introduce two independent structural interventions to isolate each degree of freedom:
  1. 1.
    The Spherical Residual Stream (§3.2): Constrains magnitude growth to inject a geometric inductive bias aligned with Fourier representations.
  2. 2.
    Uniform Attention Ablation (§3.3): Ablates data-dependent routing to test if complex attention is merely a memorization artifact, reducing aggregation to a theoretically optimal uniform sum.
Finally, to distinguish task-specific geometric alignment from generic stabilization effects, we introduce symmetric group composition ($S_{5}$) as a negative control (§3.5).
3.1 The Standard Residual Stream (Baselines)
In standard transformer architectures, the residual stream $h$ serves as a cumulative communication channel across layers. In the absence of strict topological constraints, information may be encoded in both the direction (angle) and the magnitude of the state vector. For a given layer $l$, the unconstrained forward pass is:
$h_{l}^{(mid)}$ $=h_{l-1}+\text{Attention}(h_{l-1})$ (1)
$h_{l}$ $=h_{l}^{(mid)}+\text{MLP}(h_{l}^{(mid)})$ (2)
Modern implementations commonly include normalization mechanisms such as LayerNorm or RMSNorm to stabilize optimization. However, these methods do not impose a strict bound on residual magnitude. Although activations are normalized statistically, learnable affine parameters—particularly the scaling parameter $\gamma$—allow the network to rescale representations after normalization. As a result, magnitude remains an adjustable representational degree of freedom.
Recent theoretical and empirical work suggests that overparameterized models initially converge to high-frequency or memorization-based solutions before discovering structured representations Gillman et al. (2025); Zheng et al. (2024). Motivated by these findings, we hypothesize that retained magnitude flexibility in the residual stream provides a high-variance pathway for encoding training-specific noise, fitting the training data without aligning with the task’s underlying periodic structure.
3.2 The Spherical Residual Stream and Fully Bounded Topology (Intervention A)
To examine the role of magnitude variation, we introduce the Spherical Residual Stream. We define a projection operator $\Pi_{S}$ that applies strict $L_{2}$ normalization across the feature dimension:
$\Pi_{S}(x)=\frac{x}{\max(\|x\|_{2},\epsilon)}$
where $\epsilon=10^{-8}$ ensures numerical stability. This operator is applied to the residual stream prior to each sub-layer and immediately after each residual addition. A single transformer block is thus modified as:
$h_{in}$ $=\Pi_{S}(h_{l-1})$ (3)
$h_{mid}$ $=\Pi_{S}(h_{in}+\text{Attention}(h_{in}))$ (4)
$h_{l}$ $=\Pi_{S}(h_{mid}+\text{MLP}(h_{mid}))$ (5)
This mechanism replaces standard LayerNorm, fixing the global residual norm to unity at each projection step.
Crucially, this spherical constraint acts as a direct geometric inductive bias by eliminating radial degrees of freedom in the residual stream. In high-dimensional spaces, unconstrained magnitude allows representations to spread outward and partition space using norm-based separation. As demonstrated by Zhong et al. (2023), such expansive magnitude freedom enables the construction of highly fragmented, disjoint decision regions characteristic of the memorization-heavy “Pizza” algorithm.
By projecting activations onto a fixed-norm hypersphere, we remove magnitude as a representational axis, restricting the model to encode information purely through angular relationships. While the hypersphere remains $(d_{\text{model}}-1)$-dimensional, eliminating radial variability substantially constrains how the space can be partitioned. This restriction reduces the network’s ability to form piecewise magnitude-based memorization basins and instead biases learning toward continuous angular structure.
In modular addition, this angular structure aligns naturally with 1D Fourier modes (the “Clock” algorithm). Empirically, we observe that enforcing this constraint from initialization dramatically shortens the delayed generalization phase, consistent with the hypothesis that removing magnitude-based representational freedom accelerates convergence toward structured solutions.
Fully Bounded Topology.
While the spherical residual stream provides the necessary geometry, leaving the output layer unconstrained allows the final linear unembedding matrix ($W_{\text{unembed}}$) to scale arbitrarily. Training under cross-entropy loss can drive increasing logit magnitudes to artificially lower the loss. This behavior—known as Naïve Loss Minimization—leads to numerical instability and Softmax Collapse, which can severely delay or crash generalization Prieto et al. (2025).
To stabilize the geometry and prevent Softmax Collapse, we additionally apply $\Pi_{S}$ to the unembedding weights and compute logits via scaled cosine similarity:
$\text{Logits}=\tau\left(\Pi_{S}(h_{\text{final}})\,\Pi_{S}(W_{\text{unembed}})\right)$
Since cosine similarity lies in [1,1][-1,1], logit magnitudes are strictly bounded by the temperature parameter $\tau$ (set to $10.0$ in our experiments). By constraining both the internal residual geometry and the output logit scale, this fully bounded spherical topology allows the network to stably lock into the Fourier solution without relying on weight decay to control magnitude growth.
3.3 Uniform Attention Routing (Intervention B)
[…]
This implies that for strictly commutative operations, complex query-key routing is an unnecessary degree of freedom. We hypothesize that this excess flexibility permits the network to construct asymmetric representations that memorize specific $(a,b)$ token pairs, acting as an independent driver of delayed generalization.
To formally test whether learned routing is computationally required to solve the task, or if it merely traps the model in a memorization basin, we introduce a secondary, independent architectural intervention: the Uniform Attention Ablation. In this configuration, we completely ablate the data-dependent query-key routing by overriding the pre-softmax attention scores to zero:
$\text{Scores}=\frac{QK^{T}}{\sqrt{d_{\text{head}}}}\rightarrow\mathbf{0}$ (6)
Passing this zeroed matrix through the standard Softmax operator forces a perfectly uniform attention distribution across the sequence. For a sequence containing three tokens (e.g., $a$, $b$, and ==), the attention weights become fixed at [1/3, 1/3, 1/3][1/3,\,1/3,\,1/3]. This intervention reduces the attention layer to a data-independent Continuous Bag-of-Words (CBOW) aggregator, structurally enforcing permutation invariance and perfectly matching the theoretical bag-of-tokens requirement Huang and Li (2025).
3.4 Task 1: The Cyclic Group $Z_{p}$ (Geometric Alignment and Attention)
Prior mechanistic interpretability studies show that modular addition ($\mathbb{Z}_{p}$) is frequently implemented via discrete Fourier representations, mapping integer inputs to the roots of unity on a complex circle Nanda et al. (2023). In our experiments, we set $p=113$. We generate the exhaustive set of all $p\times p$ pairs, formatting the input sequences as [a, b, equals_token]. To induce the delayed generalization characteristic of grokking, we train on a constrained fraction of the data, using a random $30\%$ split for training ($\sim 3,830$ samples) and the remaining $70\%$ for testing ($\sim 8,939$ samples):
$E(x)_{k}=\exp\left(i\frac{2\pi kx}{p}\right)=\cos\left(\frac{2\pi kx}{p}\right)+i\sin\left(\frac{2\pi kx}{p}\right)$ (7)
In standard unconstrained networks, the model must simultaneously learn to align phases, suppress magnitude variations, and route tokens appropriately to compute these trigonometric interactions. We expect that by isolating and structurally removing the degrees of freedom that allow for non-Fourier memorization—either by mathematically bounding the geometry (Intervention A) or structurally enforcing commutative token aggregation (Intervention B)—the model will bypass the prolonged memorization phase entirely.
3.5 Task 2: The Symmetric Group $S_{5}$ (Negative Control)
To isolate the effect of task-specific geometric alignment from generic optimization stabilization, we employ the composition of the symmetric group $S_{5}$ as a negative control. The $S_{5}$ composition task was originally introduced as a benchmark for delayed generalization by Power et al. (2022). The dataset consists of the pairwise composition of the 120 permutations of five elements ($a\circ b=c$). Similar to the modular addition task, we format the input sequences as [a, b, equals_token] and enforce a strict data scarcity regime by utilizing a random $30\%$ split for training ($4,320$ samples) and $70\%$ for testing ($10,080$ samples).
Unlike modular addition, $S_{5}$ is strictly non-commutative ($a\circ b\neq b\circ a$). Prior interpretability analyses of transformers trained on this task indicate that successful models construct inherently non-abelian representation structures—whether through higher-dimensional irreducible representations Chughtai et al. (2023) or subgroup coset circuits Stander et al. (2024)—rather than relying on low-dimensional circular manifolds Nanda et al. (2023). In contrast to the cyclic group $\mathbb{Z}_{p}$, whose representations reduce to 1D Fourier modes lying on a circle, empirical analyses suggest that successful solutions to $S_{5}$ composition typically involve architectures capable of supporting higher-dimensional, non-commutative structure.
Importantly, we do not claim that a spherical constraint is fundamentally incompatible with permutation composition. Rather, our hypothesis is empirical: if the acceleration observed in modular addition arises from geometric alignment between the spherical constraint and the task’s intrinsic symmetry structure, then imposing the same constraint on $S_{5}$—whose learned solutions are typically higher-dimensional and non-commutative—should not yield similar acceleration.
Specifically, if the spherical topology functions merely as a generic regularizer or stabilizer, it should reduce the grokking delay across tasks. Conversely, if acceleration depends on task-specific alignment, then constraining representations to a fixed-norm spherical manifold may restrict the model’s ability to construct the higher-dimensional structures typically observed in successful $S_{5}$ solutions, leading to delayed or absent generalization under the same training regime.
Demonstrating such a differential outcome would provide further support for our central hypothesis: that eliminating or shortening the memorization phase depends on aligning architectural degrees of freedom with the mathematical symmetries most naturally exploited by the task.
4 Empirical Results
In this section, we evaluate the hypothesis that excess representational degrees of freedom—specifically, residual magnitude flexibility and data-dependent attention routing—contribute to delayed generalization in modular arithmetic tasks. We evaluate our two structural interventions across standard baselines (LayerNorm, RMSNorm) and our bounded configurations (Spherical Norm with $\lambda=1.0$, and Fully Bounded with $\lambda=0.0$). All reported metrics are aggregated across 10 independent random seeds per configuration.
Experimental Setup.
Across all configurations, we employ a shallow Transformer architecture with $d_{\text{model}}=128$, $4$ attention heads, and an MLP hidden dimension of $512$. Models are trained using full-batch gradient descent with the AdamW optimizer. Depending on the intervention, we evaluate learning rates of $1\times 10^{-4}$ and $6\times 10^{-4}$ alongside weight decay values of $\lambda=1.0$ or $\lambda=0.0$. Comprehensive hyperparameter configurations, including the specific adjustments for the $S_{5}$ task, are detailed in Appendix A.
4.1 Intervention A: Accelerating Grokking via Topological Constraint
We first evaluate generalization dynamics under standard learned attention at a stable learning rate ($10^{-4}$). As shown in Table 1, both LayerNorm and RMSNorm baselines exhibit characteristic grokking behavior: models rapidly achieve perfect training accuracy but require an extended optimization plateau before test accuracy rises, with mean generalization epochs of 54,160 and 51,240 respectively.
Inspection of the overall learning trajectories (visualized in Figure 1) reveals the stark contrast between unconstrained and topologically bounded models. During the extended optimization plateau, the baseline models exhibit pronounced oscillations in test accuracy—intermittent gradient spikes and transient drops in generalization before eventual stabilization—consistent with the “slingshot effect” reported in prior work Thilak et al. (2022); Prieto et al. (2025). To avoid masking this behavior, gradients were not clipped during training. By contrast, the spherical and fully bounded architectures (clustered at the extreme left of Figure 1) entirely bypass these chaotic optimization dynamics, locking into the generalizing solution smoothly and immediately.
(a) Macro Dynamics (100,000 Epochs)
(b) Early Dynamics (First 5,000 Epochs)
Figure 1: Test accuracy across architectural configurations. (a) Standard normalizations exhibit the classic, delayed grokking profile over 100,000 epochs. (b) A zoomed-in view of the first 5,000 epochs reveals that the topologically bounded models bypass this delay entirely, exhibiting immediate and stable convergence while the baselines remain trapped at chance accuracy.
Table 1: Grokking onset epoch across 10 random seeds. Spherical configurations substantially reduce convergence time at both standard and elevated learning rates.
Learning Rate Architecture Mean Grok Epoch Std Dev Min Max Failures Peak Acc.
$1\times 10^{-4}$ LayerNorm (Baseline) 54,160 13,490 32,800 71,600 0 / 10 100%
RMSNorm (Baseline) 51,240 11,200 38,800 74,600 0 / 10 100%
Spherical Norm ($\lambda=1.0$) 2,480 464 1,800 3,200 0 / 10 100%
Fully Bounded ($\lambda=0.0$) 2,100 316 1,800 2,600 0 / 10 100%
$6\times 10^{-4}$ LayerNorm (Baseline) 7,800 1,095 6,000 9,400 0 / 10 100%
RMSNorm (Baseline) 7,300 925 6,000 9,200 0 / 10 100%
Spherical Norm ($\lambda=1.0$) 820 199 600 1,200 0 / 10 100%
Fully Bounded ($\lambda=0.0$) 700 194 400 1,000 0 / 10 100%
In contrast to the baselines, the spherical configurations exhibit substantially earlier generalization. At $10^{-4}$, the Fully Bounded topology reaches 100% test accuracy in a mean of 2,100 epochs—over an order of magnitude faster than the statistical normalization baselines. At a higher learning rate ($6\times 10^{-4}$), baselines require approximately 7,500 epochs on average, while the Fully Bounded configuration generalizes in roughly 700 epochs.
Empirical Validation of the Bounded Topology.
As theorized in Section 3.2, standard unconstrained models rely on explicit regularization to counteract Naïve Loss Minimization and prevent Softmax Collapse Prieto et al. (2025). When the spherical constraint was applied exclusively to the residual stream while retaining standard weight decay ($\lambda=1.0$), the model exhibited pronounced optimization instability. In this configuration, logit confidence must be achieved through scaling in the unconstrained unembedding layer, while weight decay simultaneously penalizes such growth. Empirically, this interaction coincided with large gradient spikes and variance across seeds.
In contrast, the fully bounded spherical topology—employing a normalized unembedding layer and zero weight decay ($\lambda=0.0$)—eliminates unbounded logit scaling as a degree of freedom. Under this configuration, gradient norms remained smooth, and the model reached 100% generalization rapidly. This indicates that under the fully bounded topology, stable optimization can emerge without reliance on weight decay regularization.
4.2 Intervention B: Bypassing Grokking via Uniform Attention Ablation
Independent of magnitude constraints, we tested whether the capacity for data-dependent query-key routing contributes to the observed grokking delay. Based on theoretical proofs that modular addition requires only a uniform “bag-of-tokens” aggregation Huang and Li (2025), we applied the Uniform Attention Ablation (forcing attention weights to a static $[1/3,1/3,1/3]$). Models were trained for 20,000 epochs to definitively evaluate whether they could escape the memorization basin without adaptive routing.
Table 2: Peak test accuracy under Uniform Attention Ablation (Zero-Attention) across 10 random seeds. Ablating data-dependent routing allows normalized models to completely bypass the grokking delay.
Architecture (Uniform Attention) Mean Peak Acc. Max Peak Acc. Success Rate (100% Acc)
LayerNorm Baseline (WD=1.0) 100.00% 100.00% 10 / 10
Spherical Norm (WD=1.0) 97.65% 100.00% 6 / 10
Fully Bounded Sphere (WD=0.0) 100.00% 100.00% 10 / 10
As shown in Table 2, the standard LayerNorm baseline equipped with uniform attention achieves 100% peak test accuracy across all 10 independent seeds, entirely bypassing the prolonged grokking plateau. The Fully Bounded spherical topology likewise maintains 100% success under uniform attention.
Beyond peak accuracy, examining the training dynamics under uniform attention reveals the impact of topological constraints on optimization stability. As illustrated in Figure 2, both the standard LayerNorm baseline (Figure 2a) and the Fully Bounded topology (Figure 2c) exhibit an immediate, concurrent rise in training and test accuracy, bypassing the extended memorization plateau observed in other settings. In contrast, when the spherical constraint is applied while retaining weight decay (Figure 2b), the optimization trajectory becomes markedly unstable. The tension between the model increasing logit magnitudes to reduce cross-entropy loss and weight decay penalizing that growth appears to introduce substantial optimization friction, resulting in delayed generalization.
These observations suggest that removing magnitude-based representational freedom—via a fully bounded manifold combined with zero weight decay—can promote stable, phase-free convergence in this task setting.
More broadly, the results indicate that complex, data-dependent routing is not strictly necessary for this commutative operation. By reducing the attention mechanism to a Continuous Bag-of-Words-style uniform aggregator, we eliminate symmetry-breaking routing pathways, which in this setting leads to immediate generalization.
(a) LayerNorm (Baseline)
(b) Spherical Norm ($\lambda=1.0$)
(c) Fully Bounded ($\lambda=0.0$)
Figure 2: Training dynamics under Uniform Attention Ablation (Seed 2). (a) The LayerNorm baseline immediately generalizes when data-dependent routing is removed, bypassing grokking. (b) Imposing a spherical constraint while retaining weight decay introduces optimization instability, delaying generalization. (c) The Fully Bounded topology (zero weight decay) stabilizes the geometry, resulting in immediate and flawless generalization.
4.3 Spectral Verification of the Fourier Circuit
While test accuracy confirms successful generalization, it does not reveal how the models solve the task. Prior mechanistic interpretability work shows that transformers trained on modular addition construct a Fourier circuit, mapping discrete inputs onto a continuous 1D circular manifold Nanda et al. (2023).
To evaluate whether our topological constraints preserve or distort this structure, we performed a spectral analysis of the learned representations. Our goal is to determine whether acceleration arises from earlier construction of the canonical Fourier circuit or from an alternative representational strategy. We constructed the effective linear transformation $W_{L}=W_{U}W_{out}$ and identified the top five highest-magnitude frequencies via Fast Fourier Transform (FFT). We then conducted an ablation by restricting the logits to these top five frequencies. Across successful configurations, this ablation preserved $99\%$ test accuracy (e.g., 99.98% for LayerNorm, 99.96% for Fully Bounded), indicating that these components strictly dominate the generalization mechanism. This confirms that the accelerated models rely on the same Fourier-based solution identified in prior work, rather than an alternative shortcut mechanism.
To quantify geometric structure, we computed the Fraction of Variance Explained (FVE). For each dominant frequency $k$, we measured the proportion of variance in the MLP activations explained by the ideal 2D basis functions $\cos(\omega_{k}(a+b))$ and $\sin(\omega_{k}(a+b))$.
Baseline Circuit.
The LayerNorm baseline constructs a recognizable Fourier circuit, with dominant frequencies explaining up to $\sim 54\%$ of the variance in MLP activations. In successful runs, coherent Fourier structure is observed after the prolonged optimization plateau described in Section 4.1.
Optimization Friction (Sphere, $\lambda=1.0$).
When spherical normalization is applied to the residual stream while retaining an unconstrained unembedding and standard weight decay ($\lambda=1.0$), circuit coherence degrades substantially. The maximum FVE does not exceed $\sim 29\%$, with several dominant frequencies explaining less than 10% of activation variance. This fragmentation is consistent with the optimization instability arising from the interaction between loss-driven logit scaling and weight decay regularization.
Fully Bounded Manifold (Sphere, $\lambda=0.0$).
In contrast, the fully bounded topology (normalized unembedding, $\lambda=0.0$) exhibits strong spectral alignment. Dominant frequencies explain a substantial fraction of activation variance (e.g., FVE U: 62.55%, V: 48.43%). By removing magnitude scaling from both the residual stream and the unembedding layer, optimization converges rapidly to a geometrically coherent Fourier representation.
Bounding the residual stream alone does not reliably produce coherent Fourier structure; stronger spectral alignment is observed when magnitude scaling is eliminated throughout the full pipeline.
4.4 Task-Specific Alignment: Negative Control on $S_{5}$
To determine whether the acceleration observed in Intervention A (the Fully Bounded topology) is the result of a generic optimization stabilizer or a task-specific geometric inductive bias, we evaluated the architectures on the composition of the symmetric group $S_{5}$. As detailed in Section 3.5, $S_{5}$ is non-commutative and requires higher-dimensional representational spaces that do not reduce to the continuous 1D circular manifold governing modular addition.
Because this task is fundamentally more complex, all models failed to converge within 100,000 epochs at previous learning rates; therefore, we trained all models for up to 100,000 epochs at an elevated learning rate of $10^{-3}$. The results are summarized in Table 3.
Table 3: Grokking onset epoch for the $S_{5}$ permutation composition task. Metrics for baselines are calculated exclusively over successful seeds. While the baselines generalize on the majority of seeds, the strict spherical constraints fail to produce generalization within the 100,000-epoch training window.
Architecture Mean Grok Epoch Std Dev Min Max Failures
LayerNorm (Baseline) 39,900 11,942 24,400 58,400 2 / 10
RMSNorm (Baseline) 38,250 23,600 19,000 91,600 2 / 10
Spherical Norm ($\lambda=1.0$) Failed 10 / 10
Fully Bounded ($\lambda=0.0$) Failed 10 / 10
Crucially, while the standard baselines successfully grok $S_{5}$ (mean $\sim$40,000 epochs on successful runs), the bounded spherical topologies fail to achieve generalization on any seed within the 100,000-epoch training window. Despite reaching 100% training accuracy, the spherical models remain confined to a memorization plateau, with test accuracy remaining near chance throughout training.
This differential outcome provides evidence that the Fully Bounded topology functions as a task-specific geometric inductive bias rather than a generic optimization accelerator. If the spherical constraint merely improved optimization dynamics, comparable acceleration would be expected across tasks. Instead, we observe acceleration on modular addition but consistent failure on $S_{5}$ under the same training regime.
These results support the hypothesis that shortening delayed generalization depends on alignment between architectural degrees of freedom and the symmetry structure most naturally exploited by the task. In the commutative case ($\mathbb{Z}_{p}$), the spherical constraint aligns with the circular Fourier geometry underlying successful solutions. In contrast, for the non-commutative $S_{5}$ task—where prior analyses indicate higher-dimensional representation structure—the same constraint appears to hinder the construction of generalizing circuits in this experimental setting.
5 Discussion and Conclusion
5.1 From Post-Hoc Analysis to Predictive Intervention
Mechanistic interpretability has traditionally functioned as a post-hoc observational science: models are trained under standard architectural assumptions, and circuits are subsequently reverse-engineered to explain emergent behavior. In contrast, this work demonstrates an interventional methodology. Building on prior findings that grokking in modular addition coincides with the emergence of Fourier representations, we directly restricted the architecture’s degrees of freedom to better align with the known Fourier structure of the task prior to training.
By imposing an $L_{2}$ spherical constraint (Intervention A), we structurally aligned the residual stream with the continuous circular manifold required for Fourier features, substantially accelerating generalization (over an order of magnitude in our setting). Similarly, by ablating data-dependent routing and enforcing uniform token aggregation (Intervention B), we matched the theoretical commutative requirements of the task, eliminating the delayed memorization phase in this setting. This bidirectional result—further validated by the failure of the spherical constraint on the non-commutative $S_{5}$ task—provides strong interventional evidence that architectural topology can function not merely as a descriptive lens, but as a predictive probe of task–model alignment. Representation analysis informs architectural intervention, and architectural intervention tests mechanistic hypotheses.
5.2 Toward Task-Specific Structural Alignment
Our results do not imply the existence of a single universal architectural bias. Rather, they suggest a task-by-task research program: (1) identify the representation and routing structure that naturally emerges during successful generalization, (2) encode this structure as an architectural prior, and (3) evaluate whether the memorization-to-generalization transition is shortened or eliminated.
The negative result on $S_{5}$ serves a critical complementary role. Because the spherical constraint is not aligned with the non-abelian representation structure underlying permutation composition, it fails to produce generalization under the same training setup. This contrast reinforces the hypothesis that delayed generalization is highly sensitive to the alignment between architectural degrees of freedom and task symmetry, rather than being solely an unavoidable optimization artifact. Extending alignment to non-abelian or hierarchical symmetry structures remains an open and technically nontrivial direction for future work.
5.3 The Bitter Lesson and Modality-Specific Debugging
For domains such as natural language, the underlying symmetry structure may be heterogeneous, hierarchical, or only approximately harmonic. In such cases, hard-coding a single global geometric prior is unlikely to suffice. This perspective does not contradict Sutton’s Bitter Lesson (Sutton, 2019); rather, it reframes structural alignment as a diagnostic and interventional tool for domains where mathematical structure is known or controllable.
Transformers are increasingly deployed beyond text, including in time-series forecasting, reinforcement learning, and structured decision-making tasks. In such settings, recent architectural interventions provide suggestive empirical support for symmetry-aligned design. For example, Gillman et al. (2025) demonstrate that replacing a standard linear classification head with a topologically constrained “Fourier head” significantly improves performance on continuous decision-making tasks. Similarly, Sun et al. (2025) show that explicitly incorporating periodic and Fourier-based group attention mechanisms in the PENGUIN architecture yields strong results in long-horizon time-series forecasting.
While these works do not explicitly analyze grokking dynamics, they are consistent with the hypothesis advanced here: when architectural components are aligned with the intrinsic mathematical structure of a task, models may be able to bypass memorization-heavy regimes and more directly construct invariant representations.
Recent work has also begun extending grokking beyond synthetic algorithmic tasks into real-world reasoning domains. Abramov et al. (2025) demonstrate that augmenting sparse knowledge graphs with synthetic relational data can induce grokking-like phase transitions in multi-hop factual reasoning benchmarks. Viewed through our structural lens, such results reinforce the broader perspective that grokking reflects a transition from memorization to structured relational encoding. Our contribution complements this direction by showing that this transition need not be induced post hoc through data manipulation; when architectural degrees of freedom are restricted to match task symmetry, the delayed phase can disappear entirely.
5.4 The Role of Synthetic Tasks and Scaling Geometric Bias
While our empirical validation is conducted on synthetic algorithmic datasets, this is a deliberate methodological choice consistent with the foundational literature. The phenomenon of grokking itself was initially discovered and characterized within controlled environments. Similarly, the subsequent post-hoc recoveries of Fourier circuits, non-commutative coset structures (such as those required for $S_{5}$ permutation composition), and the recent identification of magnitude-driven Softmax Collapse were all achieved using mathematical toy tasks. These controlled settings are requisite for isolating structural phase transitions from the statistical noise of heterogeneous data. Consequently, our interventional methodology builds upon these environments to provide rigorous, mathematical validation of structural alignment.
An important open question is whether these targeted constraints can improve performance beyond controlled symmetry tasks. A natural next step is to evaluate hybrid bounded-unconstrained architectures at moderate scale on heterogeneous corpora. If structural constraints primarily benefit domains with strong mathematical or cyclic symmetry, one would expect performance gains to concentrate in reasoning or algorithmic benchmarks while leaving purely linguistic benchmarks unaffected. However, such hybrid designs risk representation entanglement: without explicit routing regularizations, the model might incorrectly force unstructured linguistic features through the constrained spherical pathway. This could degrade overall performance and obfuscate post-hoc mechanistic analysis.
5.5 Conclusion
We provide controlled interventional evidence that, in modular arithmetic tasks, grokking reflects a representational realignment process rather than solely an unavoidable optimization phase transition. When a model possesses excess architectural degrees of freedom—such as unbounded magnitude scaling or complex data-dependent routing—it can rely on memorization-heavy strategies before constructing structured representations. By isolating these degrees of freedom and restricting them to better match the intrinsic commutative and periodic symmetries of modular addition, we show that the prolonged generalization delay can be dramatically shortened or, in this experimental setting, eliminated.
More broadly, this work proposes a shift from post-hoc interpretability toward predictive structural debugging: a framework in which mechanistic representation analysis informs architectural design, and architectural design in turn provides experimental tests of interpretability hypotheses. While general-purpose AI systems will continue to rely on large, flexible architectures, task-specific structural alignment offers a principled pathway for isolating and studying the factors that influence generalization dynamics.
Code Availability
All code necessary to reproduce the experiments in this work, including model implementations, training scripts, hyperparameter configurations, and analysis utilities for spectral verification and Fourier decomposition, is publicly available at: https://github.com/AlperYildirim1/geometric-grokking
Appendix A Experimental Details and Hyperparameters
All experiments were implemented in PyTorch and executed deterministically across 10 random seeds to ensure exact reproducibility. Both tasks utilized the same core Transformer architecture, with task-specific optimization parameters as outlined below.
A.1 Core Architecture
  • Embedding / Residual Dimension ($d_{\text{model}}$): 128
  • Attention Heads: 4
  • Head Dimension ($d_{\text{head}}$): 32
  • MLP Hidden Dimension ($d_{\text{mlp}}$): 512
  • Activation Function: ReLU
  • Layers: 1
A.2 Task 1: Modular Addition ($\mathbb{Z}_{113}$)
  • Sequence Format: [token_a, token_b, equals_token]
  • Vocabulary Size: 114 (tokens 0–112, plus 113 as the operator token)
  • Dataset Split: $30\%$ Train ($\sim 3,830$ samples) / $70\%$ Test ($\sim 8,939$ samples)
  • Optimizer: AdamW ($\beta_{1}=0.9,\beta_{2}=0.999$). For the Uniform Attention Ablation experiments (Intervention B), $\beta_{2}$ was set to $0.98$.
  • Batch Size: Full-batch (all training samples processed simultaneously)
  • Learning Rate: $1\times 10^{-4}$ and $6\times 10^{-4}$
  • Weight Decay: $1.0$ (Baselines & Sphere $\lambda=1.0$) / $0.0$ (Fully Bounded Sphere $\lambda=0.0$)
  • Max Epochs: 15,000 to 100,000 (depending on configuration)
A.3 Task 2: Permutation Composition ($S_{5}$)
  • Sequence Format: [token_a, token_b, equals_token]
  • Vocabulary Size: 121 (tokens 0–119, plus 120 as the operator token)
  • Dataset Split: $30\%$ Train ($4,320$ samples) / $70\%$ Test ($10,080$ samples)
  • Optimizer: AdamW ($\beta_{1}=0.9,\beta_{2}=0.999$)
  • Batch Size: Full-batch
  • Learning Rate: $1\times 10^{-3}$
  • Weight Decay: $1.0$ (Baselines & Sphere $\lambda=1.0$) / $0.0$ (Fully Bounded Sphere $\lambda=0.0$)
  • Max Epochs: 100,000

### 10 Mar 2026 v2

@@
 Abstract
-Mechanistic interpretability has largely focused on post-hoc analysis of trained networks.
-In this work, we instead adopt an interventional approach: we test mechanistic hypotheses a priori by modifying architectural topology and observing the resulting training dynamics.
-We study grokking—the delayed generalization phenomenon observed when Transformers are trained on cyclic modular addition ($\mathbb{Z}_{p}$)—and investigate whether specific architectural degrees of freedom contribute to the prolonged memorization phase.
-We identify two independent structural factors present in standard Transformers: unbounded representational magnitude and data-dependent attention routing.
-First, we introduce a fully bounded spherical topology that enforces $L_{2}$ normalization throughout the residual stream and normalizes the unembedding matrix with a fixed temperature scale.
-This constraint removes magnitude-based degrees of freedom and reduces grokking onset time by more than $20\times$ without weight decay.
-Second, we introduce a Uniform Attention Ablation that overrides data-dependent query–key routing with a fixed uniform distribution, reducing the attention layer to a Continuous Bag-of-Words (CBOW) aggregator.
-Despite eliminating adaptive routing, LayerNorm models with uniform attention achieve 100% generalization across all seeds and bypass the grokking delay.
-To evaluate whether this acceleration reflects task-specific geometric alignment rather than a generic optimization stabilizer, we employ non-commutative $S_{5}$ permutation composition as a negative control.
-Enforcing the same spherical constraints on $S_{5}$ does not accelerate generalization, suggesting that eliminating the memorization phase depends strongly on aligning architectural priors with the intrinsic symmetries of the task.
-Together, these findings provide interventional evidence that architectural degrees of freedom substantially influence the presence and duration of grokking, suggesting a predictive architectural perspective on training dynamics.
+Mechanistic interpretability typically relies on post-hoc analysis of trained networks.
+We instead adopt an interventional approach: testing hypotheses a priori by modifying architectural topology to observe training dynamics.
+We study grokking—delayed generalization in Transformers trained on cyclic modular addition ($\mathbb{Z}_{p}$)—investigating if specific architectural degrees of freedom prolong the memorization phase.
+We identify two independent structural factors in standard Transformers: unbounded representational magnitude and data-dependent attention routing.
+First, we introduce a fully bounded spherical topology enforcing $L_{2}$ normalization throughout the residual stream and an unembedding matrix with a fixed temperature scale.
+This removes magnitude-based degrees of freedom and qualitatively alters the training dynamics: models generalize immediately, bypassing the prolonged memorization phase characteristic of grokking in standard Transformer baselines, both with and without weight decay.
+Second, a Uniform Attention Ablation overrides data-dependent query-key routing with a uniform distribution, reducing the attention layer to a Continuous Bag-of-Words (CBOW) aggregator.
+Despite removing adaptive routing, these models achieve $100\%$ generalization across all seeds and bypass the grokking delay entirely.
+To evaluate whether this acceleration is a task-specific geometric alignment rather than a generic optimization stabilizer, we use non-commutative $S_{5}$ permutation composition as a negative control.
+Enforcing spherical constraints on $S_{5}$ does not accelerate generalization.
+This suggests eliminating the memorization phase depends strongly on aligning architectural priors with the task’s intrinsic symmetries.
+Together, these findings provide interventional evidence that architectural degrees of freedom substantially influence grokking, suggesting a predictive structural perspective on training dynamics.
 Keywords: Grokking, Mechanistic Interpretability, Inductive Bias, Architectural Intervention, Transformer, Representational Geometry
@@
 This intervention removes the model’s ability to encode information in vector norm and constrains logits to a bounded cosine geometry.
-Empirically, this modification substantially accelerates generalization, reducing grokking onset from $\sim$54,160 epochs in standard LayerNorm baselines to $\sim$2,100 epochs without weight decay.
+Empirically, this modification substantially accelerates generalization, reducing grokking onset from $\sim$54,160 epochs in standard LayerNorm baselines to $\sim$2,480 epochs without weight decay.
 Path B: The Routing Degree of Freedom.
@@
 While prior solutions involve relying on massive pre-training to discover Fourier features (Zhou et al., 2024) or explicitly constraining the output layer’s geometry (Gillman et al., 2025), our interventional approach tests whether enforcing a similar circular topology directly within the core residual stream can fundamentally alter the network’s optimization trajectory.
+Accelerating Grokking via Optimization.
+Complementary to geometric and representational analyses, Lee et al.
+(2024) demonstrated that grokking can be accelerated at the optimizer level by amplifying slow-varying gradient components via low-pass filtering.
+Their GROKFAST algorithm achieves up to $50\times$ faster generalization by boosting the low-frequency signal responsible for the delayed phase transition.
+However, this approach operates within the standard memorize-then-generalize paradigm: the model still undergoes a memorization phase before transitioning to generalization.
+In contrast, our architectural interventions eliminate the memorization phase entirely—bounded spherical models generalize from initialization without ever entering the prolonged overfitting plateau.
 3 Formal Problem Setting and Methodology
@@
 $\Pi_{S}(x)=\frac{x}{\max(\|x\|_{2},\epsilon)}$
-where $\epsilon=10^{-8}$ ensures numerical stability.
+where $\epsilon=10^{-12}$ ensures numerical stability.
 This operator is applied to the residual stream prior to each sub-layer and immediately after each residual addition.
@@
 Empirically, we observe that enforcing this constraint from initialization dramatically shortens the delayed generalization phase, consistent with the hypothesis that removing magnitude-based representational freedom accelerates convergence toward structured solutions.
-Fully Bounded Topology.
-While the spherical residual stream provides the necessary geometry, leaving the output layer unconstrained allows the final linear unembedding matrix ($W_{\text{unembed}}$) to scale arbitrarily.
-Training under cross-entropy loss can drive increasing logit magnitudes to artificially lower the loss.
+Bounded Output Topology.
+The spherical residual stream alone is insufficient: if the final unembedding matrix ($W_{\text{unembed}}$) remains unconstrained, training under cross-entropy loss can drive increasing logit magnitudes to artificially lower the loss.
 This behavior—known as Naïve Loss Minimization—leads to numerical instability and Softmax Collapse, which can severely delay or crash generalization Prieto et al.
 (2025).
-To stabilize the geometry and prevent Softmax Collapse, we additionally apply $\Pi_{S}$ to the unembedding weights and compute logits via scaled cosine similarity:
+Empirically, we found that all spherical models required unembedding normalization to train stably.
+To stabilize the geometry and prevent Softmax Collapse, we apply $\Pi_{S}$ to the unembedding weights and compute logits via scaled cosine similarity:
 $\text{Logits}=\tau\left(\Pi_{S}(h_{\text{final}})\,\Pi_{S}(W_{\text{unembed}})\right)$
@@
 By constraining both the internal residual geometry and the output logit scale, this fully bounded spherical topology allows the network to stably lock into the Fourier solution without relying on weight decay to control magnitude growth.
+Fourier Initialization.
+In addition to the standard random initialization, we evaluate a variant in which the first 10 dimensions of the token embedding matrix are deterministically initialized with cosine and sine values at five key frequencies, following the Fourier structure identified in prior mechanistic analyses (Nanda et al., 2023).
+This provides the model with an explicit periodic prior from initialization.
+We refer to this variant as Bounded Sphere + Fourier Init.
 3.3 Uniform Attention Routing (Intervention B)
@@
 In this section, we evaluate the hypothesis that excess representational degrees of freedom—specifically, residual magnitude flexibility and data-dependent attention routing—contribute to delayed generalization in modular arithmetic tasks.
-We evaluate our two structural interventions across standard baselines (LayerNorm, RMSNorm) and our bounded configurations (Spherical Norm with $\lambda=1.0$, and Fully Bounded with $\lambda=0.0$).
+We evaluate our two structural interventions across standard baselines (LayerNorm, RMSNorm) and our bounded spherical topology tested with weight decay ($\lambda=1.0$) and without ($\lambda=0.0$).
 All reported metrics are aggregated across 10 independent random seeds per configuration.
@@
 RMSNorm (Baseline) 51,240 11,200 38,800 74,600 0 / 10 100%
-Spherical Norm ($\lambda=1.0$) 2,480 464 1,800 3,200 0 / 10 100%
-Fully Bounded ($\lambda=0.0$) 2,100 316 1,800 2,600 0 / 10 100%
+Bounded Sphere ($\lambda=1.0$) 2,400 353 1,800 3,000 0 / 10 100%
+Bounded Sphere ($\lambda=0.0$) 2,480 464 1,800 3,200 0 / 10 100%
+Bounded Sphere + Fourier Init ($\lambda=1.0$) 2,120 368 1,600 2,600 0 / 10 100%
+Bounded Sphere + Fourier Init ($\lambda=0.0$) 2,100 316 1,800 2,600 0 / 10 100%
 $6\times 10^{-4}$ LayerNorm (Baseline) 7,800 1,095 6,000 9,400 0 / 10 100%
 RMSNorm (Baseline) 7,300 925 6,000 9,200 0 / 10 100%
-Spherical Norm ($\lambda=1.0$) 820 199 600 1,200 0 / 10 100%
-Fully Bounded ($\lambda=0.0$) 700 194 400 1,000 0 / 10 100%
+Bounded Sphere ($\lambda=1.0$) 1,560 853 600 3,000 0 / 10 100%
+Bounded Sphere ($\lambda=0.0$) 820 199 600 1,200 0 / 10 100%
+Bounded Sphere + Fourier Init ($\lambda=1.0$) 800 163 600 1,000 0 / 10 100%
+Bounded Sphere + Fourier Init ($\lambda=0.0$) 711 203 400 1,000 0 / 10 100%
 In contrast to the baselines, the spherical configurations exhibit substantially earlier generalization.
-At $10^{-4}$, the Fully Bounded topology reaches 100% test accuracy in a mean of 2,100 epochs—over an order of magnitude faster than the statistical normalization baselines.
-At a higher learning rate ($6\times 10^{-4}$), baselines require approximately 7,500 epochs on average, while the Fully Bounded configuration generalizes in roughly 700 epochs.
+At $10^{-4}$, the Bounded Sphere ($\lambda=0.0$) reaches 100% test accuracy in a mean of 2,480 epochs—over an order of magnitude faster than the statistical normalization baselines.
+At a higher learning rate ($6\times 10^{-4}$), baselines require approximately 7,500 epochs on average, while the Bounded Sphere ($\lambda=0.0$) generalizes in roughly 820 epochs.
 Empirical Validation of the Bounded Topology.
@@
 (2025).
-When the spherical constraint was applied exclusively to the residual stream while retaining standard weight decay ($\lambda=1.0$), the model exhibited pronounced optimization instability.
-In this configuration, logit confidence must be achieved through scaling in the unconstrained unembedding layer, while weight decay simultaneously penalizes such growth.
-Empirically, this interaction coincided with large gradient spikes and variance across seeds.
-In contrast, the fully bounded spherical topology—employing a normalized unembedding layer and zero weight decay ($\lambda=0.0$)—eliminates unbounded logit scaling as a degree of freedom.
-Under this configuration, gradient norms remained smooth, and the model reached 100% generalization rapidly.
-This indicates that under the fully bounded topology, stable optimization can emerge without reliance on weight decay regularization.
+All spherical configurations employ the same bounded architecture: $L_{2}$-normalized residual stream, normalized unembedding, and fixed temperature scaling.
+The only variable is weight decay.
+At the lower learning rate ($1\times 10^{-4}$), weight decay has minimal effect—both $\lambda=1.0$ and $\lambda=0.0$ converge in $\sim$2,400–2,480 epochs.
+At the higher learning rate ($6\times 10^{-4}$), however, weight decay introduces optimization friction: the $\lambda=1.0$ configuration exhibits higher variance across seeds (std 853 vs.
+199) and slower mean convergence (1,560 vs.
+820 epochs).
+This suggests that under the bounded topology, weight decay penalizes weight magnitudes that are already geometrically constrained, introducing unnecessary regularization pressure at aggressive learning rates.
+Crucially, bounding the unembedding alone does not eliminate delayed generalization.
+When an $L_{2}$-normalized unembedding is applied to the LayerNorm baseline, the model still exhibits a prolonged memorization phase.
+This occurs because LayerNorm preserves internal radial freedom in the residual stream, allowing the network to increase residual magnitudes in order to reduce cross-entropy loss through Naïve Loss Minimization Prieto et al.
+(2025).
+While recent work proposes numerical stabilizations (e.g., StableMax) that allow models to remain numerically stable during this unconstrained scaling phase Prieto et al.
+(2025), our bounded spherical topology instead removes this internal magnitude degree of freedom at the architectural level.
+By structurally constraining the residual stream itself, the architecture bypasses the prolonged memorization phase observed in baseline models and exhibits immediate generalization.
 4.2 Intervention B: Bypassing Grokking via Uniform Attention Ablation
@@
 LayerNorm Baseline (WD=1.0) 100.00% 100.00% 10 / 10
-Spherical Norm (WD=1.0) 97.65% 100.00% 6 / 10
-Fully Bounded Sphere (WD=0.0) 100.00% 100.00% 10 / 10
+Bounded Sphere ($\lambda=1.0$) 97.65% 100.00% 6 / 10
+Bounded Sphere ($\lambda=0.0$) 100.00% 100.00% 10 / 10
 As shown in Table 2, the standard LayerNorm baseline equipped with uniform attention achieves 100% peak test accuracy across all 10 independent seeds, entirely bypassing the prolonged grokking plateau.
-The Fully Bounded spherical topology likewise maintains 100% success under uniform attention.
+The Bounded Sphere ($\lambda=0.0$) likewise maintains 100% success under uniform attention.
 Beyond peak accuracy, examining the training dynamics under uniform attention reveals the impact of topological constraints on optimization stability.
-As illustrated in Figure 2, both the standard LayerNorm baseline (Figure 2a) and the Fully Bounded topology (Figure 2c) exhibit an immediate, concurrent rise in training and test accuracy, bypassing the extended memorization plateau observed in other settings.
+As illustrated in Figure 2, both the standard LayerNorm baseline (Figure 2a) and the Bounded Sphere ($\lambda=0.0$) (Figure 2c) exhibit an immediate, concurrent rise in training and test accuracy, bypassing the extended memorization plateau observed in other settings.
 In contrast, when the spherical constraint is applied while retaining weight decay (Figure 2b), the optimization trajectory becomes markedly unstable.
@@
 (a) LayerNorm (Baseline)
-(b) Spherical Norm ($\lambda=1.0$)
-(c) Fully Bounded ($\lambda=0.0$)
+(b) Bounded Sphere ($\lambda=1.0$)
+(c) Bounded Sphere ($\lambda=0.0$)
 Figure 2: Training dynamics under Uniform Attention Ablation (Seed 2).
@@
 (b) Imposing a spherical constraint while retaining weight decay introduces optimization instability, delaying generalization.
-(c) The Fully Bounded topology (zero weight decay) stabilizes the geometry, resulting in immediate and flawless generalization.
+(c) The Bounded Sphere without weight decay ($\lambda=0.0$) stabilizes the geometry, resulting in immediate and flawless generalization.
 4.3 Spectral Verification of the Fourier Circuit
@@
 In successful runs, coherent Fourier structure is observed after the prolonged optimization plateau described in Section 4.1.
-Optimization Friction (Sphere, $\lambda=1.0$).
-When spherical normalization is applied to the residual stream while retaining an unconstrained unembedding and standard weight decay ($\lambda=1.0$), circuit coherence degrades substantially.
+Optimization Friction (Bounded Sphere, $\lambda=1.0$).
+When the bounded spherical topology is trained with standard weight decay ($\lambda=1.0$), circuit coherence degrades substantially.
 The maximum FVE does not exceed $\sim 29\%$, with several dominant frequencies explaining less than 10% of activation variance.
 This fragmentation is consistent with the optimization instability arising from the interaction between loss-driven logit scaling and weight decay regularization.
-Fully Bounded Manifold (Sphere, $\lambda=0.0$).
-In contrast, the fully bounded topology (normalized unembedding, $\lambda=0.0$) exhibits strong spectral alignment.
+Bounded Sphere ($\lambda=0.0$).
+In contrast, the same bounded topology trained without weight decay ($\lambda=0.0$) exhibits strong spectral alignment.
 Dominant frequencies explain a substantial fraction of activation variance (e.g., FVE U: 62.55%, V: 48.43%).
 By removing magnitude scaling from both the residual stream and the unembedding layer, optimization converges rapidly to a geometrically coherent Fourier representation.
-Bounding the residual stream alone does not reliably produce coherent Fourier structure; stronger spectral alignment is observed when magnitude scaling is eliminated throughout the full pipeline.
 4.4 Task-Specific Alignment: Negative Control on $S_{5}$
@@
 RMSNorm (Baseline) 38,250 23,600 19,000 91,600 2 / 10
-Spherical Norm ($\lambda=1.0$) Failed 10 / 10
-Fully Bounded ($\lambda=0.0$) Failed 10 / 10
+Bounded Sphere ($\lambda=1.0$) Failed 10 / 10
+Bounded Sphere ($\lambda=0.0$) Failed 10 / 10
 Crucially, while the standard baselines successfully grok $S_{5}$ (mean $\sim$40,000 epochs on successful runs), the bounded spherical topologies fail to achieve generalization on any seed within the 100,000-epoch training window.
 Despite reaching 100% training accuracy, the spherical models remain confined to a memorization plateau, with test accuracy remaining near chance throughout training.
-This differential outcome provides evidence that the Fully Bounded topology functions as a task-specific geometric inductive bias rather than a generic optimization accelerator.
+This differential outcome provides evidence that the bounded spherical topology functions as a task-specific geometric inductive bias rather than a generic optimization accelerator.
 If the spherical constraint merely improved optimization dynamics, comparable acceleration would be expected across tasks.
@@
 Similarly, Sun et al.
-(2025) show that explicitly incorporating periodic and Fourier-based group attention mechanisms in the PENGUIN architecture yields strong results in long-horizon time-series forecasting.
+(2025) show that explicitly incorporating a modulo-based periodic relative attention bias in the PENGUIN architecture yields strong results in long-horizon time-series forecasting by structurally aligning the attention mechanism with the data’s intrinsic cyclic symmetry.
 While these works do not explicitly analyze grokking dynamics, they are consistent with the hypothesis advanced here: when architectural components are aligned with the intrinsic mathematical structure of a task, models may be able to bypass memorization-heavy regimes and more directly construct invariant representations.
@@
 While general-purpose AI systems will continue to rely on large, flexible architectures, task-specific structural alignment offers a principled pathway for isolating and studying the factors that influence generalization dynamics.
+Acknowledgements
+Large language models were used in the preparation of this manuscript for limited assistance with language polishing, LaTeX formatting of tables and figures, and routine programming tasks.
+All experimental design, implementation, analysis, and scientific conclusions were developed and verified by the author, who takes full responsibility for the content of this work.
 Code Availability
@@
   • Learning Rate: $1\times 10^{-4}$ and $6\times 10^{-4}$
-  • Weight Decay: $1.0$ (Baselines & Sphere $\lambda=1.0$) / $0.0$ (Fully Bounded Sphere $\lambda=0.0$)
+  • Weight Decay: $1.0$ (Baselines & Sphere $\lambda=1.0$) / $0.0$ (Bounded Sphere $\lambda=0.0$)
   • Max Epochs: 15,000 to 100,000 (depending on configuration)
@@
   • Learning Rate: $1\times 10^{-3}$
-  • Weight Decay: $1.0$ (Baselines & Sphere $\lambda=1.0$) / $0.0$ (Fully Bounded Sphere $\lambda=0.0$)
+  • Weight Decay: $1.0$ (Baselines & Sphere $\lambda=1.0$) / $0.0$ (Bounded Sphere $\lambda=0.0$)
   • Max Epochs: 100,000